A 10 cm radius coil of wire is hooked up to a variable current source that produces a time-varying magnetic field of magnitude B(t)=100(t-0.02t2) near the center of the coil. A secondary 1000 turn coil of radius 1 cm is placed adjacent to the first coil. Calculate the maximum induced EMF generated in the secondary coil. Assuming that the secondary coil is made of aluminum, calculate the current induced in the secondary coil.
Answer #1

the expression B(t)=100(t-0.02t2) ...is not very cleear here....i am taking it to be B(t)=100(t-0.02t^2) and solving....

emf induced= -number of turns * area of one loop * dB/dt

V= -1000*pi*(.01)^2 *100(1-.04t) , from this expression, we can see the maximum EMF induce will be infinite when t ~infinite.

to find the current we need resistivity (which can be obtained from literature) but we also need the thickness of the secondrary coil which is not given